URGENT!!! PLEASE SOLVE! Solve the system of equations. Make sure to check your solution.
y = 1/3x + 10
x = 3y + 6

Answer by stanbon(75887)

y = (1/3)x + 10
x = 3y + 6
----
Rearrange:
y = (1/3)x + 10
y = (1/3)x - 2
----
Subtracting you would get:
0 = 8, which is a contradiction.
----
Conclusion: There is no solution.
The lines are parallel; notice that they have the same slope.
==========
Cheers,
Stan H.